Wiseman Video plans to make four annual deposits of $6,500 each to a special building fund. The fund’s assets will be invested in mortgage instruments expected to pay interest at 12% on the fund’s balance. (FV of $1PV of $1FVA of $1PVA of $1FVAD of $1 and PVAD of $1(Use appropriate factor(s) from the tables provided.)

Determine how much will be accumulated in the fund on December 31, 2024 after four years, under each of the following situations.

 

  1. The first $6,500 annual deposit is made at the end of each of the four years on December 31, 2021, and interest is compounded annually.

  2. The first $6,500 annual deposit is made at the beginning of each of the four years on December 31, 2020, and interest is compounded annually.

  3. The first $6,500 annual deposit is made at the beginning of each of the four years on December 31, 2020, and interest is compounded quarterly.

  4. The first $6,500 annual deposit is made at the beginning of each of the four years on December 31, 2020, interest is compounded annually, and interest earned is withdrawn at the end of each year.
Answer #1

1) Table FVA of $1        deposit= 6,500

N= 4yrs          I= 12%            FVA= 4.7793

Fund balance: 6500*4.7793= 31,065.45     Answer: 31,065


2) Table: FVAD of $1        deposit= 6,500

N= 4yrs      i= 12%             FVAD= 5.3528

Fund balance: 6500*5.3528= 34,793.2    Answer: 34,793


3) answer:

deposit datei=n=depositfund balance
12/31/20203%166,50010,430
12/31/20213%126,5009,267
12/31/20223%86,5008,234
12/31/20233%46,5007,315



total35,246

Quarterly= 4    i=12%/4= 3%      n=4*4yrs= 16

table: FV of $1

how I got answer:        from the table

row 1) 6,500*1.60471    (3%,16)            = 10,430.615

row 2) 6,500*1.42576    (3%,12)            = 9,267.44

row 3) 6,500*1.26677     (3%,8)               = 8,234.005

row 4) 6,500*1.12551     (3%,4)             = 7,315.815


4) Answer:

deposit amount# of payments

interest left in

 fund

fund balance
6,5004026,000

6,500*4= 26,000

answered by: anonymous